OpenStudy (anonymous):

For the function f(x)= 4x+12, which of the following expressions is equivalent to [f(x)]squared (sorry didn't know how to put squared using the laptop) a.) (4x+12)squared b.) 16(xsquared +9) c.) 16xsquared+12 d.) (4(x+3))squared e.)4(x+3)squared f.)16(x+3)squared

jimthompson5910 (jim_thompson5910):

type ^2 to mean squared examples x squared = x^2 7 squared = 7^2

jimthompson5910 (jim_thompson5910):

so you want to find [f(x)]^2

jimthompson5910 (jim_thompson5910):

f(x)= 4x+12 [f(x)]^2 = (4x+12)^2 ... square both sides now you must use the FOIL rule or a table like this one  to expand out (4x+12)^2